OpenStudy (anonymous):

Find a number whose product with 8 is the same as its sum with 49. (Hint: you may want to try guess and check or use a table.) 8 7 6 9

OpenStudy (accessdenied):

Here, I would prefer using actual Algebra. Let's call the unknown number "x." Then, we can write the problem in a more mathematical way: 8x (the product of this number and 8) x + 49 (the sum of this number and 49) They're equal, so we set them equal to each other and solve for our number: 8x = x + 49

OpenStudy (accessdenied):

From here, all we have to do is subtract off the x from both sides to get the x-terms alone: 8x - x = x + 49 - x 7x = 49 Then divide by 7: x = 7

OpenStudy (accessdenied):

Otherwise, if we really wanted to guess and check, we'd just use those two expressions: 8* a number a number + 49 and plug in each answer choice to see which one works. I prefer using the Algebra, though, since its more concise and gives a direct answer. :P

OpenStudy (anonymous):

x = that number So 8x = x+ 49 7x = 49 x = 7 So the number is 7. No offense, but guess and check in this problem is a pretty stupid thing to do when it's so easy to solve it algebraically.

OpenStudy (anonymous):

If you want to use guess and check though, since it's MC just plug in the values 8(8) = 8 + 49 64 = 57, no, so it's not 8 8(7) -we already know that works 8(6) = 6 + 49 48 = 55, no, so it's not 6 8(9) = 9+49 72 = 58, no, so it's not 9